HIGH RESOLUTION SPECTROSCOPY OF ALKOXY RADICALS IN A SUPERSONIC FREE-JET

\underline{SANDHYA GOPALAKRISHNAN}, CHRISTOPHER C. CARTER, VADIM STAKHURSKY, and TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210.

We have recently reported the vibrationally resolved spectra of the B̃ ← X̃ transition of a number of alkoxy radicals containing up to 12 carbon atoms. The rotational band contours of some of these moderate-resolution spectra, coupled with ab initio calculations, indicate that some of the vibrational bands may result from different molecular conformers. A full rotational analysis should aid in distinguishing different conformers that may exist in the jet and give us information about their structure. We have commenced a program to obtain rotationally resolved spectra of the alkoxy radicals using our high resolution (fundamental bandwidth ∼100 MHz) Ar^{+} pumped, CW, pulse amplified dye laser system. We presently have in hand a number of such high resolution spectra. Work is in progress to obtain additional high resolution spectra and to complete detailed rotational analyses of them.